OpenStudy (anonymous):

Convert 126% to a mixed number (be sure to reduce!)

OpenStudy (anonymous):

126% = 126/100 = 63/50

OpenStudy (anonymous):

1 and 26/100. Simplify it to get 1 and 13/50

OpenStudy (dan815):